### v1

# Underwater Image Enhancement via Dehazing and Color Restoration

Abstract
With the rapid development of marine engineering projects such as marine resource extraction and oceanic surveys, underwater visual imaging and analysis has become a critical technology. Unfortunately, due to the inevitable non-linear attenuation of light in underwater environments, underwater images and videos often suffer from low contrast, blurriness, and color degradation, which significantly complicate the subsequent research. Existing underwater image enhancement methods often treat the haze and color cast as a unified degradation process and disregard their independence and interdependence, which limits the performance improvement. Here, we propose a Vision Transformer (ViT)-based network (referred to as WaterFormer) to improve the underwater image quality. WaterFormer contains three major components: a dehazing block (DehazeFormer Block) to capture the self-correlated haze features and extract deep-level features, a Color Restoration Block (CRB) to capture self-correlated color cast features, and a Channel Fusion Block (CFB) to capture fusion features within the network. To ensure authenticity, a soft reconstruction layer based on the underwater imaging physics model is included. To improve the quality of the enhanced images, we introduce the Chromatic Consistency Loss and Sobel Color Loss to train the network. Comprehensive experimental results demonstrate that WaterFormer outperforms other state-of-the-art methods in enhancing underwater images.
Index Terms—  Degradation process, ViT model, Underwater image enhancement
1 Introduction
The ocean holds vast natural resources, so deep-sea exploration and development are crucial. Underwater imaging serves as a vital carrier of marine information, and the image quality greatly influences marine engineering. However, with its floating particles, light scattering, and selective light absorption, the deep ocean environment introduces noise, blur, and color distortion [1], which have detrimental effects on underwater detection, underwater visual perception, and other underwater tasks. Underwater image enhancement (UIE) aims to mitigate these issues, restore the true colors, enhance the details, and consequently increase the data reliability and usability. Therefore, research on UIE algorithms is necessary, urgent, and of great significance.
In the early days, specialized hardware [2], multiple different images [3] or special imaging methods [4] were used to enhance the quality of underwater images. Although these methods were effective in certain scenarios, their expensive hardware and complex conditions limited their widespread use. To achieve convenience and universality, increasingly many image processing and computer vision techniques are used to address the challenges of underwater imaging. These techniques can be broadly categorized into non-physical model-based methods, physical model-based methods, and data-driven methods. Non-physical model-based methods do not rely on underwater optical propagation models. They adjust the pixels of the image in the spatial or transform domain to improve the visual quality of the image [5, 6, 7]. Although non-physical model-based methods are convenient and easy to use, they have controversial fidelity due to the lack of physical priors. Moreover, under special conditions, these methods can cause more severe color distortion and artifacts. Physical model-based methods consider the process of light propagation underwater, make assumptions, and build corresponding mathematical models to enhance underwater images. Example physical model-based methods estimate the light scattering parameters [8], medium transmission parameters [9, 10]. Physical model-based methods are image restoration methods and offer higher fidelity. However, they can fail in complex underwater environments and challenging imaging conditions when the assumptions do not hold. In recent years, data-driven techniques have gained significant momentum, and more researchers have begun using deep learning to enhance underwater images [11, 12]. With their large parameter space and powerful learning capabilities, deep learning models often outperform traditional methods when ample datasets are available. However, data-driven methods have three challenges. First, most data-driven methods face a trade-off between performance and the number of parameters. Second, they separately consider features of the underwater environment or rely on a single degradation process to guide the network, which limits their UIE performance. Third, they often cannot guarantee the authenticity of the enhanced data.
Hence, we devised WaterFormer, which is a lightweight but high-performance network that considers the atomization and color cast as inherent primary degradation features to the underwater environment to restore images. Specifically, WaterFormer uses the DehazeFormer Block proposed in  [13] to achieve self-attention for the haze features and extract deep features. Using the channel-wise self-attention matrix transformation, our Color Restoration Block (CRB) performs self-attention on color cast features to help the network infer results with a more realistic color distribution. We also introduced the Channel Fusion Block (CFB), which synthesized the global information from different features and refined it to generate more effective fusion features. Considering the constraints of the underwater imaging physical model, we incorporated an underwater soft reconstruction layer at the end of the network to enhance the fidelity of the results. Furthermore, we introduced the Chromatic Consistency Loss and Sobel Color Loss to train the network in maintaining chromatic consistency, preserving fine color details, and enhancing the image quality and generalization across various datasets.
Overall, the main contributions of this work are as follows:
  • A novel end-to-end network to enhance underwater images is proposed. Our approach uses the DehazeFormer Block, successfully transforms haze features and improves the image deblurring capability. It introduces an underwater soft reconstruction layer to ensure the authenticity of the results.
  • We integrate the CRB and CFB with an encoder-decoder structure based on the global information between channels to generate a more realistic color distribution and accurately fuse features from different stages.
  • We propose the Chromatic Consistency Loss and Sobel Color Loss to improve the training process of the network and enhance its generalization ability.
2 Related Work
2.1 Traditional Methods
Traditional methods can be broadly categorized into methods based on non-physical models and those based on physical models. Non-physical-model-based methods do not involve modeling the underwater optical transmission process. Instead, they adjust the pixels of an image in the spatial domain or transform domain using the characteristic statistics and visual perceptual analysis. These methods include contrast enhancement [14], histogram equalization [15, 16], white balance [17], wavelet transform [18], fusion-based methods [19, 20], retinex-based methods [21, 22, 23], and others. For example, Hitam et al. [16] applied the mixture Contrast Limited Adaptive Histogram Equalization to underwater images in the RGB and HSV color spaces. Singh et al. [18] performed a discrete wavelet transform on images and estimated the color correction factors that were associated with specific color casts to adjust the input image pixels. Ancuti et al. [19] separately conducted color correction and contrast enhancement on underwater images and subsequently fused the two versions based on automatically calculated weights. Although most non-physical model-based methods are simple to implement and computationally efficient, they lack prior knowledge of the optical propagation, which can result in oversaturated or undersaturated color distributions.
Physical-model-based methods are commonly used in underwater image enhancement. They involve mathematically modeling the underwater degradation process, estimating the parameters of the degradation model, and ultimately recovering clear images. These methods address various aspects of underwater image degradation, such as estimating the parameters related to the light scattering [8] and medium transmission [9, 24]. For example, Hou et al. [8] used optical response functions, including Point Spread Function (PSF) and Modulation Transfer Function (MTF), and deconvolution techniques for image restoration. Ke et al. [10] considered the color, saturation, and detail information in images to construct scene depth maps and edge maps to estimate the medium transmission rates. Galdran et al. [25] adjusted the Dark Channel Prior (DCP) method by incorporating red channel priors to estimate the medium transmission map for underwater scenes. Wang et al. [24] introduced the adaptive attenuation-curve prior to estimate medium transmission parameters. Physical-model-based methods are based on prior assumptions and offer high fidelity in image restoration. However, their applicability is often limited to specific underwater degradation phenomena, so they have lower generality.
2.2 Data-driven Methods
The use of deep learning techniques in UIE tasks has been a major trend in recent years. These approaches mainly follow two paradigms: the encoding-decoding paradigm and the generative paradigm. On one hand, methods based on the encoding-decoding paradigm typically encode images into a latent space and continuously learn low-level features and deep representations of images to reconstruct the final results through a decoder. Encoding-decoding methods often use global residual connections or gate fusion mechanisms to facilitate network learning. Li et al. [26] introduced a gate fusion network called Water-Net, which enhanced images by fusing the original image with white-balanced, histogram-equalized, and gamma-corrected versions. Water-Net can adapt to different underwater environments and enhance the image. Xue et al. [27] analyzed underwater degradation factors from two perspectives: color distortions and veil effects. They developed the Multi-Branch Multi-Variable Network (MBANet) to restore underwater images. Li et al. [28] proposed an enhancement network called Ucolor, which operated in the HSV, RGB, and Lab color spaces to encode images and decodes image features with a medium transmission map. Therefore, Ucolor can learn and selectively enhance different visual representations of images.
On the other hand, Generative Adversarial Networks (GANs) are key components of the generative paradigm, and they mainly consist of a generator and a discriminator. The generator produces samples from random noise, and the discriminator distinguishes between real and generated samples. Through adversarial learning, they iteratively improve each other until the generator can generate realistic samples. Islam et al. [29] designed an adversarial network called FUnIE based on conditional generation, where the generator was constructed using the Unet [30] architecture, and the discriminator was a Markovian discriminator [31]. Cong et al. [32] proposed PUGAN, which guided the generator to produce more interpretable results by estimating medium transmission parameters. Liu et al. [33] split the network training into weak and strong supervision stages, which enhanced its performance while decreasing the dependence on paired training data.
Most studies on UIE tasks are based on Convolutional Neural Networks (CNNs) because they can effectively capture local features in images through convolutional and pooling layers. In recent years, the Visual Transformer (ViT) [34] was proposed, and it can better capture both local and global perspectives of images [35] and gain more similarity between features than CNNs. As a result, some researchers began using ViT to handle UIE tasks. For example, Peng et al. [36] used the ViT architecture to build a generative adversarial paradigm network called U-shape. Shen et al. [37] constructed a ViT network called UDAformer that included a spatial attention mechanism and a channel attention mechanism in an encoder-decoder paradigm.
Although many data-driven algorithms exhibit excellent generalization, they often grapple with a trade-off between parameter count and performance. These methods overlook the interconnected but independent nature of the haze and color distortion phenomena. Treating these features as a singular degradation process neglects the interplay between haze removal and color restoration. Furthermore, existing approaches lack tailored loss functions for underwater image enhancement (UIE), which diminishes the learning capacity of the network. Hence, we developed a lightweight, high-performance network for haze and color distortion features with two proposed loss functions to enhance the underwater image features and optimize the learning capability of the network.
3 Proposed Method
Fig. 1: Overall structure of WaterFormer, which includes the DehazeFormer Block as the core component for feature extraction and transformation in the network. We enhanced the ability of the network to aggregate local information by incorporating the FReLU activation function in the MLP structure. Additionally, our proposed CRB and CFB contribute to enhancing the color features of underwater images and integrating features from different branches. The entire WaterFormer is a three-stage improved Unet network, where the Down-Sample and Up-Sample stages are implemented using convolution and pixelshuffle, respectively.
3.1 Overall framework and hierarchical structure
The proposed WaterFormer adopts the U-shaped[30] encoder-decoder architecture as the backbone. In the encoder part, the network uses the DehazeFormer Block structure[13] with self-correlation to refine the haze features and deep features. The Color Restoration Block (CRB) was designed to enhance the color characteristics of the feature maps. In the decoder part, in addition to refined features, we used the CFB for skip connections to efficiently fuse different features at different stages. Finally, to improve the realism of the estimated image, we incorporated an underwater soft reconstruction layer into WaterFormer at the end, which innovated based on the soft reconstruction layer in [13] and replaced the original global residual learning. Specifically, the underwater image formation process in [38] is as follows:
${{U}_{\lambda}}(x)={{I}_{\lambda}}(x)\cdot{{T}_{\lambda}}(x)+{{A}_{\lambda}}\cdot(1-{{T}_{\lambda}}(x))$ (1)
where ${{U}_{\lambda}}(x)$ is the underwater image captured by the camera; ${{I}_{\lambda}}(x)$ is the clear latent image; ${{T}_{\lambda}}(x)$ is the transmission map; ${A}_{\lambda}$ is the global background light; ${\lambda}$ is the wavelength of light in the R, G, B channels. According to Eq. 1, the recovery process of the underwater image can be described as:
${{I}_{\lambda}}(x)={{U}_{\lambda}}(x)\cdot(\frac{1}{{{T}_{\lambda}}(x)}-1)+{{A}_{\lambda}}\cdot(1-\frac{1}{{{T}_{\lambda}}(x)})+{{U}_{\lambda}}(x)$ (2)
We set $K_{\lambda}(x)=\frac{1}{{{T}_{\lambda}}(x)}-1$ and $B_{\lambda}(x)={{A}_{\lambda}}\cdot(\frac{1}{{{T}_{\lambda}}(x)}-1)$ to obtain the simplified Eq. 3:
${{I}_{\lambda}}(x)={{U}_{\lambda}}(x)\cdot{{K}_{\lambda}}(x)-{{B}_{\lambda}}(x)+{{U}_{\lambda}}(x)$ (3)
This equation can be introduced as a prior into our network. Specifically, we drove the network to predict the variable $O\in{{\mathbb{R}}^{h\times w\times 6}}$ and decomposed the variable into $K_{\lambda}(x)\in{{\mathbb{R}}^{h\times w\times 3}}$ and $B_{\lambda}(x)\in{{\mathbb{R}}^{h\times w\times 3}}$. Eq. 3 is referenced to reconstruct the final ground truth estimation image.
To reduce the complexity and computational cost of the network, we removed a significant number of convolutional layers from the original DehazeFormer Block structure and only retained the convolution operations related to the self-attention and multilayer perceptron (MLP). To compensate for the limited utilization of local information due to the decrease in number of convolutional layers, we replaced the activation function in the MLP with a more locally enhanced FReLU [39] and designed WaterFormer as a three-stage Unet-based network. If there are more complex underwater environments and less demand for network speed performance, the number of encoding-decoding stages can be increased to enhance the network performance.
3.1.1 Color Restoration Block (CRB)
Due to the absorption and scattering of light in water, underwater images tend to appear bluish or greenish compared to land images, which is a channel-level per-pixel interference. Therefore, leveraging the global information among the channels for color balancing is an important task in underwater image enhancement. Inspired by the matrix transformations in  [40, 41], we designed the CRB to perform matrix transformations on the pixel-wise information among input feature channels using a cross-channel cross-covariance. This process generates an attention map that implicitly encodes the global context, facilitates interactions within the channels and promotes the color enhancement. The computational complexity of the channel-level global self-attention is linear, so the CRB can exhibit faster computational performance than spatial self-attention computations [42, 34, 35]. Fig. 2 illustrates the specific structure of the CRB.
Fig. 2: Overall structure of the CRB. Through channel-wise self-attention, WaterFormer can better self-transform and enhance color features in images.
We used the Rescale Layer Normalization (RLN) from [13] because it ensures consistency in the mean and variance between input and output feature maps. We believe that the RLN can help preserve the image brightness, restore colors, and reduce artifacts. For a given input $x\in{{\mathbb{R}}^{c\times h\times w}}$, we first applied the RLN and used the $1\times 1$ point-wise convolution and $3\times 3$ depth-wise convolution to map the RLN into the required $Q$, $K$, and $V$ values for self-attention computation. This process can be expressed by the following Eq. 4:
$Q,K,V=Flat(DConv(PConv(RLN(x))))$ (4)
Here, $PConv(\cdot)$ is a $1\times 1$ point-wise convolution; $DConv(\cdot)$ is a $3\times 3$ depth-wise convolution; $Flat(\cdot)$ is Flattening operation; $Q,K,V\in{{\mathbb{R}}^{c\times hw}}$.
During the computation of the global channel-wise self-attention, we performed a $3\times 3$ convolution operation on the $V$ values. This step helped improve the performance of the network by gathering information from the neighboring regions. The overall process of attention calculation can be described as follows:
${Atten}=(softmax(Q\cdot{{K}^{T}})/\gamma)\cdot Conv3(V)$ (5)
where $\gamma$ is the scaling coefficient for stable computation, and $Conv3(\cdot)$ is the convolution operation with a $3\times 3$ kernel size.
Finally, we restored the mean and variance (Affine) of $Atten$, added $Atten$ pixel-wise with the input $x$ and further refined the features through an MLP layer.
3.1.2 Channel Fusion Block (CFB)
Pixel-wise addition and concatenation are the most common strategies for multi-feature map fusion. However, these methods treat the information among different feature maps equally. We believe that due to the nonlinear absorption of light by water at different wavelengths, the information between channels in underwater image features is not equally important. Therefore, in our designed CFB, the information among different feature map channels is fused based on their importance. Fig. 3 shows the specific process of the CFB. For two given feature maps $x_{1}\in{{\mathbb{R}}^{c\times h\times w}}$ and $x_{2}\in{{\mathbb{R}}^{c\times h\times w}}$, we obtained the weights $\{{\alpha}_{1}$ and ${\alpha}_{2}\}$ using the global average pooling $GAP(\cdot)$, softmax function, and split operation. Then, we performed a weighted sum and refined the sum result using $PConv$, $DConv$, and residual connections. The more formal fusion process of the CFB is as follows:
$\{{{\alpha}_{1}},{{\alpha}_{2}}\}=split(softmax(GAP({{x}_{1}}),GAP({{x}_{2}})))$ (6)
${y}^{\prime}={{\alpha}_{1}}\cdot{{x}_{1}}+{{\alpha}_{2}}\cdot{{x}_{2}}$
$y={y}^{\prime}+DConv(PConv({y}^{\prime}))$
where $y\in{{\mathbb{R}}^{c\times h\times w}}$ is the fused generated feature map.
Fig. 3: Overall structure of the CFB.
3.2 Loss Function
The overall loss function of our WaterFormer contains three terms: the Reconstruction Loss ${{L}_{\ell_{1}}}$, Chromatic Consistency Loss ${{L}_{chroma}}$ and Sobel Color Loss ${{L}_{Sobel}}$.
Reconstruction Loss. For the given normalized real image ${{I}_{gt}}\in{{\mathbb{R}}^{h\times w\times c}}$ and network predicted image ${{I}_{pred}}\in{{\mathbb{R}}^{h\times w\times c}}$, ${{L}_{\ell_{1}}}$ is calculated as follows:
${{L}_{{{\ell}_{1}}}}=\frac{1}{h\times w\times c}\cdot\sum\limits_{k=1}^{c}{\sum\limits_{n=1}^{h}{\sum\limits_{m=1}^{w}{\left|{{I}_{gt}}(k,n,m)-{{I}_{pred}}(k,n,m)\right|}}}$ (7)
Chromatic Consistency Loss. To help the network learn the color differences between ${I}_{gt}$ and ${I}_{pred}$ while maintaining consistent chromaticity, we designed ${{L}_{chroma}}$ based on the chromatic similarity term proposed in [43]. First, we transformed ${I}_{gt}$ and ${I}_{pred}$ from the $RGB$ space to the $YIQ$ space using a matrix transformation:
$\left[\begin{matrix}Y\\ I\\ Q\\ \end{matrix}\right]=\left[\begin{matrix}0.299&0.587&0.114\\ 0.596&-0.274&-0.322\\ 0.211&-0.523&0.312\\ \end{matrix}\right]\left[\begin{matrix}R\\ G\\ B\\ \end{matrix}\right]$ (8)
The $I$ and $Q$ chromaticity channels of image ${I}_{gt}$ (${I}_{pred}$) are represented by ${I}_{gt}^{I}$ (${I}_{pred}^{I}$) and ${I}_{gt}^{Q}$ (${I}_{pred}^{Q}$), respectively. The chromatic consistency values ${{S}_{chroma}^{I}}$ and ${{S}_{chroma}^{Q}}$ of channels $I$ and $Q$ are calculated as follows:
${{S}_{chroma}^{I}}=\frac{\sigma(I_{gt}^{I},I_{pred}^{I})+{{c}_{1}}}{{{\sigma}^{2}}(I_{gt}^{I})+{{\sigma}^{2}}(I_{pred}^{I})+{{c}_{1}}}$ (9)
${{S}_{chroma}^{Q}}=\frac{\sigma(I_{gt}^{Q},I_{pred}^{Q})+{{c}_{2}}}{{{\sigma}^{2}}(I_{gt}^{Q})+{{\sigma}^{2}}(I_{pred}^{Q})+{{c}_{2}}}$
where $S_{chroma}^{I},S_{chroma}^{Q}\in\left(0,1\right]$, and the values closer to 1 indicate higher chromatic consistency between the two images. ${{\sigma}^{2}(\cdot)}$ is the variance calculation; ${{\sigma}(\cdot,\cdot)}$ is the covariance calculation; $c_{1}$ and $c_{2}$ are normalization constants to harmonize the chromatic consistency, which we set to 0.001 in our experiment. We defined $L_{chroma}$ as follows:
${L_{chroma}}=1-{S_{chroma}^{I}}\cdot{S_{chroma}^{Q}}$ (10)
In the experiment, to enhance the chromatic consistency of local regions, we set a sliding window with a size of $15\times 15$ and a stride of 1 to calculate the chromatic consistency loss for each local region. Finally, we obtained the total chromatic consistency loss by taking the mean of all local losses.
Sobel Color Loss. Hazing typically results in blurriness in underwater images, which manifests as the loss of edge details and a decrease in contrast. To overcome this challenge, we effectively preserved the fine color details between ${{I}_{gt}}$ and ${{I}_{pred}}$ by constraining the contour components of different color channels. Specifically, we defined the Sobel color loss in Eq. 11:
$L_{Sobel}^{x}=\frac{1}{3}\cdot\sum\limits_{\lambda}^{\{R,G,B\}}{{{L}_{{{\ell}_{1}}}}(Sobe{{l}_{x}}(I_{gt}^{\lambda}),Sobe{{l}_{x}}(I_{pred}^{\lambda}))}$ (11)
$L_{Sobel}^{y}=\frac{1}{3}\cdot\sum\limits_{\lambda}^{\{R,G,B\}}{{{L}_{{{\ell}_{1}}}}(Sobe{{l}_{y}}(I_{gt}^{\lambda}),Sobe{{l}_{y}}(I_{pred}^{\lambda}))}$
${{L}_{Sobel}}=L_{Sobel}^{x}+L_{Sobel}^{y}$
where $Sobe{{l}_{x}}(\cdot)$ and $Sobe{{l}_{y}}(\cdot)$ are the convolution operations of $Sobel$ operator in the $x$ and $y$ directions, respectively. The final loss function is defined as:
$L={{\lambda}_{1}}\cdot{{L}_{{{\ell}_{1}}}}+{{\lambda}_{2}}\cdot{{L}_{chroma}}+{{\lambda}_{3}}\cdot{{L}_{Sobel}}$ (12)
Here, ${\lambda}_{1}$, ${\lambda}_{2}$, and ${\lambda}_{3}$ are constants that balance the proportion of each loss term, which were set as 3, 1, and 3, respectively, with numerous experiments.
4 Experiments
In this section, we present the details and settings of the experiments. Then, we evaluate the performance of the proposed model and compare it with other representative methods. Finally, we conduct a series of ablation experiments to demonstrate the effectiveness of the proposed components.
4.1 Implementation Details
Our experiments were conducted on both synthetic underwater datasets [44] and real underwater dataset UIEBD [26]. In the synthetic underwater dataset, the reference images followed real color distributions, whereas the synthetic images simulated 10 different underwater degradation environments (I, IA, IB, II, and III simulated open sea underwater environments; 1, 3, 5, 7, and 9 simulated nearshore underwater environments). Each category of synthetic images had 1449 images, which were paired with reference images. We randomly selected 300 paired images from each category as the training set. The UIEBD dataset comprised 890 pairs of real underwater-reference images and 60 real underwater images without pairs. The reference images were obtained through enhancement using various traditional methods. We randomly partitioned 800 paired images as the training set. All images were resized to $256\times 256$ for training, and we augmented the dataset by applying random flips and rotations.
The experiments were conducted on tow NVIDIA GPUs (GeForce GTX 2080Ti) using the PyTorch 1.7.1 framework on an Ubuntu 16.04 system. We trained the network using the Adam optimizer with an initial learning rate of 0.001 and decayed it by a factor of $\frac{1}{2}$ every 50 epochs. We set the number of training iterations to 300 epochs and the batch size to 4.
4.2 Experiment settings
Benchmarks. In the synthetic underwater dataset, we randomly selected 30 additional paired images from each category of synthetic data as the test set, which was denoted as Test-S300. In the UIEBD dataset, we selected the remaining 90 paired images as the test set, which was denoted as Test-U90. Test-S300 and Test-U90 served as our full-reference benchmarks for the experiments. To demonstrate the generalization performance of our proposed network, we randomly selected 60 and 16 unpaired real underwater images from the UIEBD dataset and SQUID dataset [45], respectively, as the no-reference datasets. These datasets are called Test-U60 and Test-SQ16, respectively.
Compared Methods. We compared the proposed WaterFormer to other state-of-the-art underwater image enhancement methods based on deep learning: UWCNN[44], Water-Net[26], Ucolor[28], FUnIE[29], U-shape[36], UDAformer[37], and LE-Net[46]. In addition, we compared WaterFormer with the following traditional methods: Fusion-based[19], Retinex-based[21], MLLE[14], UDCP[47], RED[25], and HFM[48].
Evaluation Metrics. For Test-S300 and Test-U90, we used SSIM, PSNR, and NRMSE as the full-reference evaluation metrics for our experiments. These metrics are widely used to assess the similarity between two images due to their convenience and accuracy. Higher values of SSIM and PSNR or lower values of NRMSE indicate greater similarity between two images. For Test-U60, we used UCIQEE [49] and UIQM [50] as the no-reference evaluation metrics. Higher values of UCIQE and UIQM indicate better visual perception of the underwater images by humans. However, UCIQE and UIQM may not accurately reflect the actual visual perception in certain cases.
Test-S300 Test-U90 Overhead
Methods SSIM↑ PSNR↑ NRMSE↓ SSIM↑ PSNR↑ NRMSE↓ #Param↓ MACs↓ Time↓
Fusion-based[19] 0.651 12.65 0.685 0.872 21.80 0.188 × × 0.073s
Retinex-based[21] 0.641 14.54 0.520 0.773 17.34 0.283 × × 0.208s
MLLE[14] 0.603 12.38 0.667 0.740 16.50 0.315 × × 0.038s
UDCP[47] 0.597 13.84 0.500 0.689 13.85 0.418 × × 1.694s
RED[25] 0.659 12.66 0.674 0.833 19.03 0.244 × × 0.020s
HFM[48] 0.676 15.00 0.377 0.835 18.32 0.293 × × 0.328s
UWCNN[44] 0.640 14.62 0.456 0.818 18.10 0.265 0.041M 2.649G 0.002s
Water-Net[26] 0.746 19.25 0.322 0.905 23.17 0.153 1.091M 71.42G 0.016s
FUnIE[29] 0.769 21.34 0.231 0.814 18.94 0.239 7.020M 10.24G 0.002s
Ucolor[28] 0.813 22.23 0.235 0.897 21.23 0.180 148.8M 404.8G 0.169s
U-shape[36] 0.793 21.89 0.213 0.806 20.28 0.208 22.82M 2.983G 0.025s
UDAformer[37] 0.798 22.39 0.226 0.921 23.38 0.147 9.594M 41.59G 0.042s
LE-Net[46] 0.774 19.19 0.277 0.865 21.42 0.180 0.014M 0.690G 0.006s
WaterFormer 0.907 29.40 0.116 0.929 24.39 0.137 0.313M 7.749G 0.023s
Table 1: Quantitative comparison results for each method on the Test-S300 and Test-U90 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.3 Quantitative Comparisons
Full-reference Evaluation. Tab. 1 shows the quantitative comparison results on Test-S300 and Test-U90. The optimal results are highlighted in red, and the second-best results are highlighted in blue. WaterFormer achieved the best SSIM, PSNR, and NRMSE metrics on both full-reference test sets. Notably, for Test-S300, WaterFormer surpassed the second-best method by 9.4% in SSIM, 31.6% in PSNR, and shows a 9.7% lower NRMSE. This significant improvement indicated that previous methods were not able to effectively restore the true color distribution, detail distribution, and noise distribution of underwater images. Similarly, for Test-U90, where different methods were used to enhance the original images to obtain reference images, WaterFormer surpassed the second-best method by 0.8% in SSIM, 4.3% in PSNR, and exhibited a 1.0% lower NRMSE. This represents that our proposed network performs better in fitting underwater image enhancement compared with the previously best-performing methods.
Overhead. In Tab. 1, we included the inference speed of each method. Additionally, for deep learning methods, we measured the overall cost of the network using the total number of parameters (#Param) and multiply-accumulate operations (MACs). All measurements of the inference speed and network cost were conducted on images of size $256\times 256$. Compared to other methods, our proposed WaterFormer achieves optimal inference results while maintaining fast inference speed and relatively low network overhead.
Non-reference Evaluation. Tab. 2 shows the quantitative comparison results on Test-U60 and Test-SQ16. The best-performing and second-best results are highlighted in red and blue, respectively. Unfortunately, our proposed WaterFormer did not achieve the best or second-best results in the UIQM and UCIQE metrics on either no-reference datasets. As referenced  [28, 36] notes, while the UIQM and UCIQE metrics effectively evaluate the color, brightness, contrast, saturation, and sharpness of underwater images, they struggle to accurately measure color artifacts and shifts. In scenarios with more color artifacts and shifts, these metrics tend to yield higher scores, contradicting human perception. Our WaterFormer consistently avoids color artifacts and shifts, contributing to our lower UIQM and UCIQE scores. Hence, while we consider UIQM and UCIQE scores as indicative, they shouldn’t be the sole criteria for assessing the algorithm performance.
Test-U60 Test-SQ16
Methond UIQM↑ UCIQE↑ UIQM↑ UCIQE↑
Input 2.042 0.478 0.813 0.393
Fusion-based[19] 2.601 0.621 1.629 0.589
Retinex-based[21] 2.677 0.582 2.336 0.597
MLLE[14] 1.967 0.593 2.340 0.558
UDCP[47] 1.639 0.554 1.067 0.599
RED[25] 2.330 0.559 1.362 0.560
HFM[48] 2.298 0.585 2.273 0.607
UWCNN[44] 2.411 0.487 2.018 0.404
Water-Net[26] 2.694 0.598 2.562 0.567
FUnIE[29] 2.927 0.560 2.599 0.491
Ucolor[28] 2.695 0.564 2.383 0.508
U-shape[36] 2.812 0.528 2.184 0.493
UDAformer[37] 2.807 0.576 2.487 0.519
LE-Net[46] 2.549 0.561 2.297 0.512
WaterFormer 2.546 0.594 2.372 0.543
Table 2: Quantitative comparison results for each method on the Test-U60 and Test-SQ16 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.4 Qualitative Comparisons
Full-reference Evaluation. Fig. 4 and Fig. 5 show the qualitative comparison results on Test-S300 and Test-U90, respectively. For the Test-S300 dataset, the images enhanced by Fusion-based and RED methods had a reddish bias and failed to recover the true image details from the haze. The Retinex-based and UWCNN methods reduced the saturation of the enhanced images and even distorted the true color tones. Water-Net and Ucolor approximated the real colors of the synthesized images, but they still introduced significant artifacts. In contrast, WaterFormer did not generate artifacts and closely restored the brightness, color tones, and details in each region of the recovered images. Regarding the Test-U90 dataset, apart from Water-Net and WaterFormer, other methods failed to recover the desired colors, introduced artifacts, or could not effectively dehaze the images. Our proposed WaterFormer could better restore the colors and details and produced more accurate results than Water-Net.
Non-reference Evaluation. Fig. 6 and Fig. 7 show the qualitative comparison results on Test-U60 and Test-SQ16 , respectively. For the Test-U60 dataset, none of the methods satisfactorily performed. For example, the Fusion-based method generated artificial colors, the UDCP method produced low-brightness results, the UWCNN method shows minimal enhancement effect, the Water-Net method had limited effect on the shark image in the third row, the FUnIE method tended to produce color artifacts, and the U-shape method yielded low contrast results. Meanwhile, the images enhanced by WaterFormer achieved satisfying brightness, contrast, and color. Similarly, for the Test-SQ16 dataset, the Fusion-based, UDCP, UWCNN, and FUnIE methods failed to effectively restore the image colors. Water-Net and FUnIE introduced color shifts and artifacts, whereas the U-shape method had disappointing contrast and edge details in the enhanced images. In contrast, the images enhanced by WaterFormer restored as many colors and details as possible but exhibited minimal color shifts and artifacts.
[…]
Fig. 4: Qualitative Comparison Results on Test-S300. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
[…]
Fig. 5: Qualitative Comparison Results on Test-U90. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
[…]
Fig. 6: Qualitative Comparison Results on Test-U60. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
[…]
Fig. 7: Qualitative Comparison Results on Test-SQ16. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
[…]
Fig. 8: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-S300, respectively. (a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
[…]
Fig. 9: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-U90, respectively. (a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
4.5 Other Comparisons
To further demonstrate that our proposed method has better color correction and broader applications than other methods, qualitative analyses were conducted on Test-S300 and Test-U90. Each image underwent histogram display, Canny edge detection, and saliency object detection. The saliency object detection method was referenced from [51], and Fig. 8 and Fig. 9 show the corresponding results. For Test-S300 and Test-U90, with the exception of WaterFormer, all methods failed to adequately reproduce the color distribution, gradient distribution, and salient feature distribution of the true/reference images. For example, the UDAformer generated color artifacts on Test-S300, which incurred inaccurate edge detection results. Meanwhile, when applied to Test-U90, MLLE uniformly processed all channels, which introduced gradient noise and diminished the salient features of the image. In summary, WaterFormer exhibited outstanding performance in the UIE task and proves to be the most beneficial in related applications.
4.6 Ablation Study
The proposed WaterFormer incorporates three major components: the CRB, CFB, and loss of combining ${{L}_{chroma}}$ and ${{L}_{Sobel}}$. To demonstrate the effectiveness of each component, an ablation study was performed on the baseline framework with various combinations of these components. Tab. 3 shows the statistical results on the Test-S300 dataset, where the full model ($V_{5}$) outperformed other variants in terms of SSIM and PSNR scores.
Variants CRB CFB ${{L}_{chroma}}$ ${{L}_{Sobel}}$ SSIM↑ PSNR↑
Base w/o w/o w/o w/o 0.872 26.64
${{V}_{1}}$ w/o w/o w/o 0.876 27.09
${{V}_{2}}$ w/o w/o 0.877 27.26
${{V}_{3}}$ w/o 0.891 28.47
${{V}_{4}}$ w/o 0.901 28.56
${{V}_{5}}$ 0.907 29.40
Table 3: Quantitative analysis of the WaterFormer ablation experiments. The red font indicates the best performance.
Fig. 10 displays the visual results of various variants of WaterFormer. The Base model has basic UIE capabilities but retained significant shortcomings in addressing the color degradation and edge artifacts. The $V_{1}$ model, which incorporates the CRB module, could more delicately restore colors. However, due to insufficient feature fusion, the results of the $V_{1}$ model remained lacking. Therefore, by introducing the CFB module, the predictions of the $V_{2}$ model more closely aligned with the true color distribution. The addition of ${{L}_{chroma}}$ in the $V_{3}$ model further enhanced the consistency between inference results and chromaticity of the real images. However, $V_{1}$, $V_{2}$, and $V_{3}$ continued to exhibit noticeable edge artifacts, and this problem was addressed by the inclusion of ${{L}_{Sobel}}$, as demonstrated in the results of the $V_{4}$ inference. Finally, by combining all components, our WaterFormer ($V_{5}$) yielded the closest results to the real images.
[…]
Fig. 10: Qualitative analysis of the WaterFormer ablation experiments. (a) Input, (b) Base, (c) $V_{1}$, (d) $V_{2}$, (e) $V_{3}$, (f) $V_{4}$, (g) $V_{5}$, (h) GT.
In addition, we will demonstrate the effectiveness of other improvements on WaterFormer in Tab. 4, including replacing the ReLU with the FReLU [39] in the MLP structure and replacing the global residual reconstruction layer with the soft reconstruction layer that was proposed in [13] and optimized according to Eq. 3. References [13, 52] also proposed the SK Fusion structure to fuse features from different stages. Unfortunately, we found that the SK Fusion structure was not suitable for feature fusion in underwater images, possibly because SK Fusion does not effectively utilize the global information among different feature channels in underwater images. We present the ablation experiments on SK Fusion in Tab. 4.
Variants SSIM↑ PSNR↑
WaterFormer 0.907 29.40
FReLU [39]→ReLU 0.905 29.40
UW Soft Recon→Recon 0.904 29.00
UW Soft Recon→Soft Recon [13] 0.904 28.93
CFB→SK fuison [13, 52] 0.903 28.88
Table 4: Quantitative analysis of the ablation experiments on other improvements in WaterFormer. The red font indicates the best performance.
5 Conclusion
In this work, we introduced WaterFormer, a versatile method to enhance the quality of underwater images. To tackle challenges such as haziness and color shifts, WaterFormer leverages the Vision Transformer architecture and comprises three key components: the DehazeFormer Block, Color Restoration Block (CRB), and Channel Fusion Block (CFB). The DehazeFormer Block transforms hazy features and extract deep-level features. The CRB enhances colors through channel-level transpose matrices. The CFB facilitates feature fusion based on the importance of global information across different feature channels. We also incorporated the Chromatic Consistency Loss and Sobel Color Loss to ensure color fidelity and preserve fine details. The quantitative evaluations, qualitative evaluations, and ablation experiments show that our proposed network performs significantly better than various methods.
Nevertheless, this study has several limitations. First, our work requires a substantial amount of paired data for training, which poses a significant challenge for practical applications. Second, we did not consider the effects of noise such as environmental noise and camera noise during the image processing. Large-scale datasets that cover and fully consider complex types of image noise will be essential to train a more robust and generalizable model.
In our subsequent work, we will consider incorporating a denoising module into the network to reduce the impact of noise and enhance the robustness of the image correction results. In addition, we will use a contrastive learning strategy to reduce the network dependency on data and improve the generalization performance. The denoising process will be integrated into the end-to-end network as a module instead of a separate preprocessing step. We will also explore the processing or analysis based on enhanced results, such as the underwater image classification and measures of image representability. We believe that the analysis of restored underwater images can provide incremental assistance in future underwater visual-imaging tasks.

### v2

@@
 Abstract
-With the rapid development of marine engineering projects such as marine resource extraction and oceanic surveys, underwater visual imaging and analysis has become a critical technology.
-Unfortunately, due to the inevitable non-linear attenuation of light in underwater environments, underwater images and videos often suffer from low contrast, blurriness, and color degradation, which significantly complicate the subsequent research.
-Existing underwater image enhancement methods often treat the haze and color cast as a unified degradation process and disregard their independence and interdependence, which limits the performance improvement.
-Here, we propose a Vision Transformer (ViT)-based network (referred to as WaterFormer) to improve the underwater image quality.
-WaterFormer contains three major components: a dehazing block (DehazeFormer Block) to capture the self-correlated haze features and extract deep-level features, a Color Restoration Block (CRB) to capture self-correlated color cast features, and a Channel Fusion Block (CFB) to capture fusion features within the network.
+Underwater visual imaging is crucial for marine engineering, but it suffers from low contrast, blurriness, and color degradation, which hinders downstream analysis.
+Existing underwater image enhancement methods often treat the haze and color cast as a unified degradation process, neglecting their inherent independence while overlooking their synergistic relationship.
+To overcome this limitation, we propose a Vision Transformer (ViT)-based network (referred to as WaterFormer) to improve underwater image quality.
+WaterFormer contains three major components: a dehazing block (DehazeFormer Block) to capture the self-correlated haze features and extract deep-level features, a Color Restoration Block (CRB) to capture self-correlated color cast features, and a Channel Fusion Block (CFB) that dynamically integrates these decoupled features to achieve comprehensive enhancement.
 To ensure authenticity, a soft reconstruction layer based on the underwater imaging physics model is included.
-To improve the quality of the enhanced images, we introduce the Chromatic Consistency Loss and Sobel Color Loss to train the network.
+Further, a Chromatic Consistency Loss and Sobel Color Loss are designed to respectively preserve color fidelity and enhance structural details during network training.
 Comprehensive experimental results demonstrate that WaterFormer outperforms other state-of-the-art methods in enhancing underwater images.
@@
 The ocean holds vast natural resources, so deep-sea exploration and development are crucial.
-Underwater imaging serves as a vital carrier of marine information, and the image quality greatly influences marine engineering.
-However, with its floating particles, light scattering, and selective light absorption, the deep ocean environment introduces noise, blur, and color distortion [1], which have detrimental effects on underwater detection, underwater visual perception, and other underwater tasks.
-Underwater image enhancement (UIE) aims to mitigate these issues, restore the true colors, enhance the details, and consequently increase the data reliability and usability.
+Underwater imaging serves as a vital carrier of marine information and image quality greatly influences marine engineering.
+However, with its floating particles, light scattering and selective light absorption, the deep ocean environment introduces noise, blur, and color distortion [1], which have detrimental effects on underwater detection, underwater visual perception, and other underwater tasks [2, 3].
+Underwater Image Enhancement (UIE) aims to mitigate these issues, restore the true colors, enhance the details, and consequently increase the reliability and usability of the data.
 Therefore, research on UIE algorithms is necessary, urgent, and of great significance.
-In the early days, specialized hardware [2], multiple different images [3] or special imaging methods [4] were used to enhance the quality of underwater images.
+In the early days, specialized hardware [4], multiple different images [5] or special imaging methods [6] were used to improve the quality of underwater images.
 Although these methods were effective in certain scenarios, their expensive hardware and complex conditions limited their widespread use.
@@
 Non-physical model-based methods do not rely on underwater optical propagation models.
-They adjust the pixels of the image in the spatial or transform domain to improve the visual quality of the image [5, 6, 7].
-Although non-physical model-based methods are convenient and easy to use, they have controversial fidelity due to the lack of physical priors.
-Moreover, under special conditions, these methods can cause more severe color distortion and artifacts.
-Physical model-based methods consider the process of light propagation underwater, make assumptions, and build corresponding mathematical models to enhance underwater images.
-Example physical model-based methods estimate the light scattering parameters [8], medium transmission parameters [9, 10].
-Physical model-based methods are image restoration methods and offer higher fidelity.
-However, they can fail in complex underwater environments and challenging imaging conditions when the assumptions do not hold.
-In recent years, data-driven techniques have gained significant momentum, and more researchers have begun using deep learning to enhance underwater images [11, 12].
+They adjust in the spatial or transform domain to improve the visual quality of the image [7, 8].
+Non-physical model-based methods are user-friendly but may exacerbate color distortion or introduce unnatural artifacts due to their reliance on statistical/empirical adjustments, which ignore the physics of wavelength-selective light attenuation underwater.
+Moreover, most approaches decouple dehazing and color correction into separate stages, causing iterative error accumulation: information lost in earlier stages cannot be retrieved later.
+Physical model-based methods enhance underwater images by analyzing the principles of light propagation, formulating assumptions, and constructing mathematical models to simulate underwater processes.
+For example, these approaches often estimate critical parameters such as light scattering coefficients [9] and medium transmission rates [10, 11].
+Using the analogous modeling frameworks shared between underwater imaging and atmospheric haze imaging, these methods achieve simultaneous color restoration and dehazing with high fidelity.
+However, their effectiveness diminishes in complex underwater environments due to the overdependence of simplified assumptions, such as uniform illumination and static water conditions, which fail to account for real-world challenges such as artificial directional lighting, dynamic turbulence, and suspended particulate matter.
+Additionally, the interdependent nature of color correction and dehazing introduces mutual performance degradation when ideal conditions for either task (e.g., accurate parameter estimation or stable scene properties) are not met.
+In recent years, data-driven techniques have gained significant momentum and more researchers have begun to use deep learning to enhance underwater images [12, 13].
 With their large parameter space and powerful learning capabilities, deep learning models often outperform traditional methods when ample datasets are available.
-However, data-driven methods have three challenges.
+However, data-driven methods face three challenges.
 First, most data-driven methods face a trade-off between performance and the number of parameters.
-Second, they separately consider features of the underwater environment or rely on a single degradation process to guide the network, which limits their UIE performance.
-Third, they often cannot guarantee the authenticity of the enhanced data.
-Hence, we devised WaterFormer, which is a lightweight but high-performance network that considers the atomization and color cast as inherent primary degradation features to the underwater environment to restore images.
+Second, most of them only focus on deep feature extraction or color-based guidance, neglecting haze features’ importance, which limits their UIE performance in dehazing and edge detail retention.
+Third, they often cannot guarantee the authenticity of the enhanced data due to the lack of physical model guidance.
+Hence, we devised WaterFormer, which is a lightweight but high-performance network that considers the haze and color cast as inherent primary degradation features of the underwater environment to restore images.
 Specifically, WaterFormer uses the DehazeFormer Block proposed in [14] to achieve self-attention for the haze features and extract deep features.
@@
 Considering the constraints of the underwater imaging physical model, we incorporated an underwater soft reconstruction layer at the end of the network to enhance the fidelity of the results.
-Furthermore, we introduced the Chromatic Consistency Loss and Sobel Color Loss to train the network in maintaining chromatic consistency, preserving fine color details, and enhancing the image quality and generalization across various datasets.
+Furthermore, we introduced the Chromatic Consistency Loss and Sobel Color Loss to train the network to maintain chromatic consistency, preserve fine color details, and enhance image quality and generalization across various datasets.
 Overall, the main contributions of this work are as follows:
-  • A novel end-to-end network to enhance underwater images is proposed.
-Our approach uses the DehazeFormer Block, successfully transforms haze features and improves the image deblurring capability.
+  • A novel end-to-end network is proposed to enhance underwater images.
+Our approach uses the DehazeFormer Block, successfully transforms haze features, and improves the image deblurring capability.
 It introduces an underwater soft reconstruction layer to ensure the authenticity of the results.
@@
 Traditional methods can be broadly categorized into methods based on non-physical models and those based on physical models.
-Non-physical-model-based methods do not involve modeling the underwater optical transmission process.
+Non-physical model-based methods do not involve modeling the underwater optical transmission process.
 Instead, they adjust the pixels of an image in the spatial domain or transform domain using the characteristic statistics and visual perceptual analysis.
-These methods include contrast enhancement [14], histogram equalization [15, 16], white balance [17], wavelet transform [18], fusion-based methods [19, 20], retinex-based methods [21, 22, 23], and others.
+These methods include contrast enhancement [15], histogram equalization [16], wavelet transform [17], fusion-based methods [18], retinex-based methods [19], and others.
 For example, Hitam et al.
@@
 Although most non-physical model-based methods are simple to implement and computationally efficient, they lack prior knowledge of the optical propagation, which can result in oversaturated or undersaturated color distributions.
-Physical-model-based methods are commonly used in underwater image enhancement.
-They involve mathematically modeling the underwater degradation process, estimating the parameters of the degradation model, and ultimately recovering clear images.
+Physical model-based methods are commonly used in UIE.
+They involve mathematical modeling of the underwater degradation process, estimating the parameters of the degradation model, and ultimately recovering clear images.
 These methods address various aspects of underwater image degradation, such as estimating the parameters related to the light scattering [9] and medium transmission [10, 20].
@@
 [21] adjusted the Dark Channel Prior (DCP) method by incorporating red channel priors to estimate the medium transmission map for underwater scenes.
-Wang et al.
-[24] introduced the adaptive attenuation-curve prior to estimate medium transmission parameters.
-Physical-model-based methods are based on prior assumptions and offer high fidelity in image restoration.
+Physical model-based methods are based on prior assumptions and offer high fidelity in image restoration.
 However, their applicability is often limited to specific underwater degradation phenomena, so they have lower generality.
@@
 These approaches mainly follow two paradigms: the encoding-decoding paradigm and the generative paradigm.
-On one hand, methods based on the encoding-decoding paradigm typically encode images into a latent space and continuously learn low-level features and deep representations of images to reconstruct the final results through a decoder.
+On the one hand, methods based on the encoding-decoding paradigm typically encode images into a latent space and continuously learn low-level features and deep representations of images to reconstruct the final results through a decoder.
 Encoding-decoding methods often use global residual connections or gate fusion mechanisms to facilitate network learning.
@@
 Li et al.
-[28] proposed an enhancement network called Ucolor, which operated in the HSV, RGB, and Lab color spaces to encode images and decodes image features with a medium transmission map.
+[23] proposed an enhancement network called Ucolor, which operated in the HSV, RGB, and Lab color spaces to encode images and decode image features with a medium transmission map.
 Therefore, Ucolor can learn and selectively enhance different visual representations of images.
-On the other hand, Generative Adversarial Networks (GANs) are key components of the generative paradigm, and they mainly consist of a generator and a discriminator.
-The generator produces samples from random noise, and the discriminator distinguishes between real and generated samples.
-Through adversarial learning, they iteratively improve each other until the generator can generate realistic samples.
+On the other hand, Generative Adversarial Networks (GANs) and Diffusion Models constitute two pivotal methodologies within the generative paradigm.
+GANs comprise two components: a generator that synthesizes samples from random noise, and a discriminator that differentiates authentic samples from synthetic ones.
+Through adversarial optimization, both components co-evolve until the generator attains photorealistic synthesis capability.
+Diffusion models involve a forward diffusion process and a reverse generation stage.
+During the forward process, incremental Gaussian noise is systematically introduced to an image until its statistical properties converge to a standard normal distribution, concurrently training a Unet [24] to estimate the noise profile.
+The reverse phase leverages the trained Unet to progressively refine the corrupted latent representation into a coherent image through iterative denoising operations.
 Islam et al.
-[29] designed an adversarial network called FUnIE based on conditional generation, where the generator was constructed using the Unet [30] architecture, and the discriminator was a Markovian discriminator [31].
+[25] designed an adversarial network called FUnIE based on conditional generation, where the generator was constructed using Unet architecture and the discriminator was a Markovian discriminator.
 Cong et al.
 [26] proposed PUGAN, which guided the generator to produce more interpretable results by estimating medium transmission parameters.
-Liu et al.
-[33] split the network training into weak and strong supervision stages, which enhanced its performance while decreasing the dependence on paired training data.
+Tang et al.
+[27] developed a diffusion model-based UIE framework, employing piecewise sampling and evolutionary algorithm search to optimize generation efficiency.
+Zhao et al.
+[28] innovatively integrated frequency domain representations into the diffusion architecture, where a Wavelet-based Fourier interaction network and Frequency Residual Diffusion Adjustment Module collectively demonstrated the critical role of spectral characteristics in UIE tasks.
 Most studies on UIE tasks are based on Convolutional Neural Networks (CNNs) because they can effectively capture local features in images through convolutional and pooling layers.
-In recent years, the Visual Transformer (ViT) [34] was proposed, and it can better capture both local and global perspectives of images [35] and gain more similarity between features than CNNs.
+In recent years, the ViT [29] was proposed, and it can better capture both local and global perspectives of images and gain more similarity between features than CNNs.
 As a result, some researchers began using ViT to handle UIE tasks.
@@
 Treating these features as a singular degradation process neglects the interplay between haze removal and color restoration.
-Furthermore, existing approaches lack tailored loss functions for underwater image enhancement (UIE), which diminishes the learning capacity of the network.
+Furthermore, existing approaches lack tailored loss functions for UIE, which decreases the learning capacity of the network.
 Hence, we developed a lightweight, high-performance network for haze and color distortion features with two proposed loss functions to enhance the underwater image features and optimize the learning capability of the network.
@@
 1: Overall structure of WaterFormer, which includes the DehazeFormer Block as the core component for feature extraction and transformation in the network.
-We enhanced the ability of the network to aggregate local information by incorporating the FReLU activation function in the MLP structure.
-Additionally, our proposed CRB and CFB contribute to enhancing the color features of underwater images and integrating features from different branches.
-The entire WaterFormer is a three-stage improved Unet network, where the Down-Sample and Up-Sample stages are implemented using convolution and pixelshuffle, respectively.
+It’s ability to aggregate local information is enhanced through the incorporation of the FReLU in the MLP.
+Additionally, the proposed CRB and CFB are employed to improve the color features of underwater images and facilitate the integration of features across different branches.
+WaterFormer is built as a three-stage, enhanced Unet network, with Down-Sample and Up-Sample stages implemented using convolution and pixelshuffle, respectively.
+Input/output images are from the EUVP dataset[25].
+Block Type Num. of Blocks Embedding Dims Num. of Heads MLP Ratio Act. Function of MLP
+DehazeFormer Block [2, 4, 2] [24, 48, 24] [3, 6, 3] [3, 4, 3] FReLU
+Color Restoration Block [2, 2] [24, 48] [3, 6] [3, 4] FReLU
+Table 1: Detailed structure of the Transformer block.
 3.1 Overall framework and hierarchical structure
 The proposed WaterFormer adopts the U-shaped[24] encoder-decoder architecture as the backbone.
-In the encoder part, the network uses the DehazeFormer Block structure[13] with self-correlation to refine the haze features and deep features.
+In the encoder part, the network uses the DehazeFormer Block[14] with self-correlation to refine the haze features and deep features.
 The Color Restoration Block (CRB) was designed to enhance the color characteristics of the feature maps.
@@
 If there are more complex underwater environments and less demand for network speed performance, the number of encoding-decoding stages can be increased to enhance the network performance.
+Fig.
+1 and Tab.
+1 present the detailed structure of WaterFormer.
 3.1.1 Color Restoration Block (CRB)
 Due to the absorption and scattering of light in water, underwater images tend to appear bluish or greenish compared to land images, which is a channel-level per-pixel interference.
-Therefore, leveraging the global information among the channels for color balancing is an important task in underwater image enhancement.
+Therefore, leveraging the global information among the channels for color balancing is an important task in UIE.
 Inspired by the matrix transformations in  [34, 35], we designed the CRB to perform matrix transformations on the pixel-wise information among input feature channels using a cross-channel cross-covariance.
 This process generates an attention map that implicitly encodes the global context, facilitates interactions within the channels and promotes the color enhancement.
-The computational complexity of the channel-level global self-attention is linear, so the CRB can exhibit faster computational performance than spatial self-attention computations [42, 34, 35].
+The computational complexity of the channel-level global self-attention is linear, so the CRB can exhibit faster computational performance than spatial self-attention computations [29].
 Fig.
@@
 $Q,K,V=Flat(DConv(PConv(RLN(x))))$ (4)
-Here, $PConv(\cdot)$ is a $1\times 1$ point-wise convolution; $DConv(\cdot)$ is a $3\times 3$ depth-wise convolution; $Flat(\cdot)$ is Flattening operation; $Q,K,V\in{{\mathbb{R}}^{c\times hw}}$.
+Here, $PConv(\cdot)$ is a $1\times 1$ point-wise convolution; $DConv(\cdot)$ is a $3\times 3$ depth-wise convolution; $Flat(\cdot)$ is a flattening operation; $Q,K,V\in{{\mathbb{R}}^{c\times hw}}$.
 During the computation of the global channel-wise self-attention, we performed a $3\times 3$ convolution operation on the $V$ values.
@@
 Pixel-wise addition and concatenation are the most common strategies for multi-feature map fusion.
-However, these methods treat the information among different feature maps equally.
+However, these methods treat information among different feature maps equally.
 We believe that due to the nonlinear absorption of light by water at different wavelengths, the information between channels in underwater image features is not equally important.
@@
 ${L_{chroma}}=1-{S_{chroma}^{I}}\cdot{S_{chroma}^{Q}}$ (10)
-In the experiment, to enhance the chromatic consistency of local regions, we set a sliding window with a size of $15\times 15$ and a stride of 1 to calculate the chromatic consistency loss for each local region.
+In the experiment, to enhance the chromatic consistency of local regions, we set a sliding window with a size of $15\times 15$ (see Tab.
+7) and a stride of 1 to calculate the chromatic consistency loss for each local region.
 Finally, we obtained the total chromatic consistency loss by taking the mean of all local losses.
@@
 $L={{\lambda}_{1}}\cdot{{L}_{{{\ell}_{1}}}}+{{\lambda}_{2}}\cdot{{L}_{chroma}}+{{\lambda}_{3}}\cdot{{L}_{Sobel}}$ (12)
-Here, ${\lambda}_{1}$, ${\lambda}_{2}$, and ${\lambda}_{3}$ are constants that balance the proportion of each loss term, which were set as 3, 1, and 3, respectively, with numerous experiments.
+Here, ${\lambda}_{1}$, ${\lambda}_{2}$, and ${\lambda}_{3}$ are constants that balance the proportion of each loss term, set to 1, 1, and 2, respectively, based on numerous experiments (see Tab. 8).
 4 Experiments
@@
 We randomly selected 300 paired images from each category as the training set.
-The UIEBD dataset comprised 890 pairs of real underwater-reference images and 60 real underwater images without pairs.
+The UIEBD dataset includes 890 paired underwater real images and reference images, as well as 60 unpaired underwater real images.
 The reference images were obtained through enhancement using various traditional methods.
@@
 The experiments were conducted on tow NVIDIA GPUs (GeForce GTX 2080Ti) using the PyTorch 1.7.1 framework on an Ubuntu 16.04 system.
-We trained the network using the Adam optimizer with an initial learning rate of 0.001 and decayed it by a factor of $\frac{1}{2}$ every 50 epochs.
-We set the number of training iterations to 300 epochs and the batch size to 4.
+We adopt the AdamW optimizer with hyperparameters ${\beta}_{1}=0.9$ and ${\beta}_{2}=0.999$, initializing the learning rate at $1\times 10^{-3}$.
+A cosine annealing scheduler (CosineAnnealingLR) progressively adjusts the learning rate from $1\times 10^{-3}$ to $1\times 10^{-7}$ over 400 training epochs, with a batch size of 4.
 4.2 Experiment settings
 Benchmarks.
-In the synthetic underwater dataset, we randomly selected 30 additional paired images from each category of synthetic data as the test set, which was denoted as Test-S300.
+In the synthetic underwater dataset, we randomly selected 30 additional paired images from each synthetic data category as the test set, which was denoted Test-S300.
 In the UIEBD dataset, we selected the remaining 90 paired images as the test set, which was denoted as Test-U90.
@@
 Compared Methods.
-We compared the proposed WaterFormer to other state-of-the-art underwater image enhancement methods based on deep learning: UWCNN[44], Water-Net[26], Ucolor[28], FUnIE[29], U-shape[36], UDAformer[37], and LE-Net[46].
+We compared the proposed WaterFormer to other state-of-the-art UIE methods based on deep learning: UWCNN[12], Water-Net[13], Ucolor[23], FUnIE[25], U-shape[30], UDAformer[31], GuidedHybSensUIR[38], DM-water[27], and WF-Diff[28].
 In addition, we compared WaterFormer with the following traditional methods: Fusion-based[18], Retinex-based[19], MLLE[15], UDCP[39], RED[21], and HFM[40].
@@
 Higher values of SSIM and PSNR or lower values of NRMSE indicate greater similarity between two images.
-For Test-U60, we used UCIQEE [49] and UIQM [50] as the no-reference evaluation metrics.
+For Test-U60, we used UCIQEE [41] and UIQM [42] as no-reference evaluation metrics.
 Higher values of UCIQE and UIQM indicate better visual perception of the underwater images by humans.
-However, UCIQE and UIQM may not accurately reflect the actual visual perception in certain cases.
-Test-S300 Test-U90 Overhead
-Methods SSIM↑ PSNR↑ NRMSE↓ SSIM↑ PSNR↑ NRMSE↓ #Param↓ MACs↓ Time↓
-Fusion-based[19] 0.651 12.65 0.685 0.872 21.80 0.188 × × 0.073s
-Retinex-based[21] 0.641 14.54 0.520 0.773 17.34 0.283 × × 0.208s
-MLLE[14] 0.603 12.38 0.667 0.740 16.50 0.315 × × 0.038s
-UDCP[47] 0.597 13.84 0.500 0.689 13.85 0.418 × × 1.694s
-RED[25] 0.659 12.66 0.674 0.833 19.03 0.244 × × 0.020s
-HFM[48] 0.676 15.00 0.377 0.835 18.32 0.293 × × 0.328s
-UWCNN[44] 0.640 14.62 0.456 0.818 18.10 0.265 0.041M 2.649G 0.002s
-Water-Net[26] 0.746 19.25 0.322 0.905 23.17 0.153 1.091M 71.42G 0.016s
-FUnIE[29] 0.769 21.34 0.231 0.814 18.94 0.239 7.020M 10.24G 0.002s
-Ucolor[28] 0.813 22.23 0.235 0.897 21.23 0.180 148.8M 404.8G 0.169s
-U-shape[36] 0.793 21.89 0.213 0.806 20.28 0.208 22.82M 2.983G 0.025s
-UDAformer[37] 0.798 22.39 0.226 0.921 23.38 0.147 9.594M 41.59G 0.042s
-LE-Net[46] 0.774 19.19 0.277 0.865 21.42 0.180 0.014M 0.690G 0.006s
-WaterFormer 0.907 29.40 0.116 0.929 24.39 0.137 0.313M 7.749G 0.023s
-Table 1: Quantitative comparison results for each method on the Test-S300 and Test-U90 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
-4.3 Quantitative Comparisons
+However, UCIQE and UIQM may not accurately reflect actual visual perception in certain cases.
+Methods Test-S300 Test-U90 Overhead
+SSIM↑ PSNR↑ NRMSE↓ SSIM↑ PSNR↑ NRMSE↓ #Param(M)↓ MACs(G)↓ Time(s)↓
+Fusion-based[18] 0.651 12.65 0.685 0.872 21.80 0.188 × × 0.073
+Retinex-based[19] 0.641 14.54 0.520 0.773 17.34 0.283 × × 0.208
+MLLE[15] 0.603 12.38 0.667 0.740 16.50 0.315 × × 0.038
+UDCP[39] 0.597 13.84 0.500 0.689 13.85 0.418 × × 1.694
+RED[21] 0.659 12.66 0.674 0.833 19.03 0.244 × × 0.020
+HFM[40] 0.676 15.00 0.377 0.835 18.32 0.293 × × 0.328
+UWCNN[12] 0.640 14.62 0.456 0.818 18.10 0.265 0.041 2.649 0.002
+Water-Net[13] 0.746 19.25 0.322 0.905 23.17 0.153 1.091 71.42 0.016
+FUnIE[25] 0.769 21.34 0.231 0.814 18.94 0.239 7.020 10.24 0.002
+Ucolor[23] 0.813 22.23 0.235 0.897 21.23 0.180 148.8 404.8 0.169
+[…]
+UDAformer[31] 0.798 22.39 0.226 0.921 23.38 0.147 9.594 41.59 0.042
+[…]
+WF-Diff[28] 0.827 22.41 0.234 0.924 23.82 0.144 18.48 122.76 0.465
+[…]
+[…]
+Table 2: Quantitative comparison results for each method on the Test-S300 and Test-U90 datasets. Bold and underline indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
+4.3 Results and Analysis
 Full-reference Evaluation.
@@
 2 shows the quantitative comparison results on Test-S300 and Test-U90.
-The optimal results are highlighted in red, and the second-best results are highlighted in blue.
-WaterFormer achieved the best SSIM, PSNR, and NRMSE metrics on both full-reference test sets.
-Notably, for Test-S300, WaterFormer surpassed the second-best method by 9.4% in SSIM, 31.6% in PSNR, and shows a 9.7% lower NRMSE.
+The optimal results are presented in bold, while the second-best results are marked with an underline.
+WaterFormer achieved excellent SSIM, PSNR, and NRMSE scores on both full-reference test sets.
+What is noteworthy is that, for the Test-S300 dataset, WaterFormer exceeded the second-best method by 1.66% in SSIM, 13.3% in PSNR, and achieved a 19.8% lower NRMSE.
 This significant improvement indicated that previous methods were not able to effectively restore the true color distribution, detail distribution, and noise distribution of underwater images.
-Similarly, for Test-U90, where different methods were used to enhance the original images to obtain reference images, WaterFormer surpassed the second-best method by 0.8% in SSIM, 4.3% in PSNR, and exhibited a 1.0% lower NRMSE.
-This represents that our proposed network performs better in fitting underwater image enhancement compared with the previously best-performing methods.
-Overhead.
-In Tab.
-1, we included the inference speed of each method.
-Additionally, for deep learning methods, we measured the overall cost of the network using the total number of parameters (#Param) and multiply-accumulate operations (MACs).
-All measurements of the inference speed and network cost were conducted on images of size $256\times 256$.
-Compared to other methods, our proposed WaterFormer achieves optimal inference results while maintaining fast inference speed and relatively low network overhead.
+For the Test-U90 dataset, WaterFormer achieved the best performance in SSIM, surpassing the second-best method, DM-water, by 0.22%.
+However, in terms of PSNR and NRMSE, WaterFormer ranked second, being 2.65% lower and 0.76% higher than DM-water, respectively.
+SSIM considers the consistency of brightness, contrast, and structural information, indicating that WaterFormer focused more on results aligned with human visual perception.
+In contrast, PSNR and NRMSE are more sensitive to pixel-level noise, suggesting a greater emphasis on noise reduction capability.
+Despite slight differences in their inferred results, the evaluation metrics were very close, demonstrating WaterFormer’s exceptional ability in image enhancement and fitting.
+[…]
+Fig. 4: Qualitative Comparison Results on Test-S300. (a) Input, (b) Fusion-based[18], (c) Retinex-based[19], (d) Water-Net[13], (e) Ucolor[23], (f) WF-Diff[28], (g) WaterFormer, (h) GT.
+[…]
+Fig. 5: Qualitative Comparison Results on Test-U90. (a) Input, (b) Fusion-based[18], (c) Retinex-based[19], (d) Water-Net[13], (e) Ucolor[23], (f) WF-Diff[28], (g) WaterFormer, (h) GT.
+Fig.
+4 and Fig.
+5 show the qualitative comparison results on Test-S300 and Test-U90, respectively.
+For the Test-S300 dataset, images enhanced using Fusion-based and Retinex-based methods exhibited noticeable reddish and blackish color biases, failing to recover the true details obscured by haze.
+Although the Water-Net, Ucolor, and WF-Diff methods succeeded in roughly restoring the original colors of the synthesized images, they still fell short of recovering the full saturation and tonal distribution.
+Furthermore, they introduced color artifacts that were absent in the original images.
+In contrast, WaterFormer did not produce such artifacts and successfully restored brightness, color tones, and details across different regions of the enhanced images.
+Regarding the Test-U90 dataset, Fusion-based and Retinex-based enhancements still resulted in red and black color biases, along with prominent red edge artifacts.
+The ability of Ucolor to recover colors was relatively limited, and the enhanced images lacked sufficient tone and saturation.
+However, the results from Water-Net, WF-Diff, and WaterFormer were closer to the reference images, demonstrating their effectiveness in accurately restoring both color and fine details from real underwater scenes.
 Non-reference Evaluation.
@@
 3 shows the quantitative comparison results on Test-U60 and Test-SQ16.
-The best-performing and second-best results are highlighted in red and blue, respectively.
+The best and second best results are presented in bold and underlined, respectively.
 Unfortunately, our proposed WaterFormer did not achieve the best or second-best results in the UIQM and UCIQE metrics on either no-reference datasets.
-As referenced  [28, 36] notes, while the UIQM and UCIQE metrics effectively evaluate the color, brightness, contrast, saturation, and sharpness of underwater images, they struggle to accurately measure color artifacts and shifts.
+As noted in  [23, 30], while the UIQM and UCIQE metrics effectively evaluate the color, brightness, contrast, saturation, and sharpness of underwater images, they struggle to accurately measure color artifacts and shifts.
 In scenarios with more color artifacts and shifts, these metrics tend to yield higher scores, contradicting human perception.
 Our WaterFormer consistently avoids color artifacts and shifts, contributing to our lower UIQM and UCIQE scores.
-Hence, while we consider UIQM and UCIQE scores as indicative, they shouldn’t be the sole criteria for assessing the algorithm performance.
-Test-U60 Test-SQ16
-Methond UIQM↑ UCIQE↑ UIQM↑ UCIQE↑
+Hence, while we consider the UIQM and UCIQE scores as indicative, they should not be the sole criteria for assessing the algorithm performance.
+Methods Test-U60 Test-SQ16
+UIQM↑ UCIQE↑ UIQM↑ UCIQE↑
 Input 2.042 0.478 0.813 0.393
@@
 MLLE[15] 1.967 0.593 2.340 0.558
-UDCP[47] 1.639 0.554 1.067 0.599
+[…]
 RED[21] 2.330 0.559 1.362 0.560
@@
 UWCNN[12] 2.411 0.487 2.018 0.404
-Water-Net[26] 2.694 0.598 2.562 0.567
+[…]
 FUnIE[25] 2.927 0.560 2.599 0.491
 Ucolor[23] 2.695 0.564 2.383 0.508
-U-shape[36] 2.812 0.528 2.184 0.493
+[…]
 UDAformer[31] 2.807 0.576 2.487 0.519
-LE-Net[46] 2.549 0.561 2.297 0.512
-WaterFormer 2.546 0.594 2.372 0.543
-Table 2: Quantitative comparison results for each method on the Test-U60 and Test-SQ16 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
-4.4 Qualitative Comparisons
-Full-reference Evaluation.
-Fig.
-4 and Fig.
-5 show the qualitative comparison results on Test-S300 and Test-U90, respectively.
-For the Test-S300 dataset, the images enhanced by Fusion-based and RED methods had a reddish bias and failed to recover the true image details from the haze.
-The Retinex-based and UWCNN methods reduced the saturation of the enhanced images and even distorted the true color tones.
-Water-Net and Ucolor approximated the real colors of the synthesized images, but they still introduced significant artifacts.
-In contrast, WaterFormer did not generate artifacts and closely restored the brightness, color tones, and details in each region of the recovered images.
-Regarding the Test-U90 dataset, apart from Water-Net and WaterFormer, other methods failed to recover the desired colors, introduced artifacts, or could not effectively dehaze the images.
-Our proposed WaterFormer could better restore the colors and details and produced more accurate results than Water-Net.
-Non-reference Evaluation.
+DM-water[27] 2.704 0.575 2.212 0.525
+WF-Diff[28] 2.785 0.576 2.336 0.520
+GuidedHybSensUIR[38] 2.757 0.584 2.440 0.556
+WaterFormer 2.594 0.581 2.395 0.551
+Table 3: Quantitative comparison results for each method on the Test-U60 and Test-SQ16 datasets. Bold and underline indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
+[…]
+Fig. 6: Qualitative Comparison Results on Test-U60. (a) Input, (b) UDCP[39], (c) MLLE[15], (d) FUnIE[25], (e) U-shape[30], (f) DM-water[27], (g) GuidedHybSensUIR[38], (h) WaterFormer.
+[…]
+Fig. 7: Qualitative Comparison Results on Test-SQ16. (a) Input, (b) UDCP[39], (c) MLLE[15], (d) FUnIE[25], (e) U-shape[30], (f) DM-water[27], (g) GuidedHybSensUIR[38], (h) WaterFormer.
 Fig.
@@
 7 show the qualitative comparison results on Test-U60 and Test-SQ16, respectively.
-For the Test-U60 dataset, none of the methods satisfactorily performed.
-For example, the Fusion-based method generated artificial colors, the UDCP method produced low-brightness results, the UWCNN method shows minimal enhancement effect, the Water-Net method had limited effect on the shark image in the third row, the FUnIE method tended to produce color artifacts, and the U-shape method yielded low contrast results.
-Meanwhile, the images enhanced by WaterFormer achieved satisfying brightness, contrast, and color.
-Similarly, for the Test-SQ16 dataset, the Fusion-based, UDCP, UWCNN, and FUnIE methods failed to effectively restore the image colors.
-Water-Net and FUnIE introduced color shifts and artifacts, whereas the U-shape method had disappointing contrast and edge details in the enhanced images.
-In contrast, the images enhanced by WaterFormer restored as many colors and details as possible but exhibited minimal color shifts and artifacts.
-[…]
-Fig. 4: Qualitative Comparison Results on Test-S300. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
-[…]
-Fig. 5: Qualitative Comparison Results on Test-U90. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
-[…]
-Fig. 6: Qualitative Comparison Results on Test-U60. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
-[…]
-Fig. 7: Qualitative Comparison Results on Test-SQ16. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
+For the Test-U60 dataset, none of the methods produced satisfactory results.
+For example, while the UDCP method performed color correction, its capability was relatively weak, and the resulting images suffered from insufficient brightness.
+The MLLE method, on the other hand, produced incorrect color recovery results, such as a green submarine body (second row) and a purple ocean (third row).
+The FUnIE method introduced color artifacts that were inaccurate.
+All of the U-shape, DM-water, and GuidedHybSensUIR methods produced enhanced images that lacked adequate contrast and sharpness.
+Unlike the above methods, the images restored by WaterFormer exhibited satisfactory brightness, contrast, sharpness, and color.
+Similarly, for the Test-SQ16 dataset, the UDCP method again produced images with insufficient brightness.
+The MLLE and FUnIE methods struggled with recovering the correct colors and introduced severe color artifacts.
+The U-shape, DM-water and GuidedHybSensUIR methods also failed to restore the contrast and sharpness of the images adequately.
+In comparison, the images generated by WaterFormer restored more colors and details as much as possible, without introducing any additional color shifts or artifacts.
+Overhead.
+In Tab.
+2, we included the inference speed of each method.
+Furthermore, for deep learning methods, we measured the total cost of the network using the total number of parameters (#Param) and multiply-accumulate operations (MACs).
+All inference speed and network cost measurements were conducted on images of size $256\times 256$.
+Compared to other methods, our proposed WaterFormer achieves optimal inference results while maintaining fast inference speed and relatively low network overhead.
 […]
@@
 8: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-S300, respectively.
-(a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
+(a) Input, (b) MLLE[15], (c) HFM[40], (d) UDAformer[31], (e) DM-water[27], (f) GuidedHybSensUIR[38], (g)WaterFormer, (h) GT.
 […]
@@
 9: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-U90, respectively.
-(a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
+(a) Input, (b) MLLE[15], (c) HFM[40], (d) UDAformer[31], (e) DM-water[27], (f) GuidedHybSensUIR[38], (g)WaterFormer, (h) GT.
 4.4 Other Comparisons
@@
 9 show the corresponding results.
-For Test-S300 and Test-U90, with the exception of WaterFormer, all methods failed to adequately reproduce the color distribution, gradient distribution, and salient feature distribution of the true/reference images.
-For example, the UDAformer generated color artifacts on Test-S300, which incurred inaccurate edge detection results.
-Meanwhile, when applied to Test-U90, MLLE uniformly processed all channels, which introduced gradient noise and diminished the salient features of the image.
+For Test-S300 and Test-U90, MLLE, HFM, and UDAformer failed to accurately reproduce the color, gradient, and saliency features of the reference images.
+For example, UDAformer introduced color artifacts in Test-S300, leading to inaccurate edge detection results.
+MLLE applied uniform processing across all channels in Test-U90, introducing gradient noise and reducing saliency features.
+Conversely, DM-water, GuidedHybSensUIR, and WaterFormer restored the information more accurately.
+Upon closer inspection, WaterFormer provided the most precise results.
+For instance, WaterFormer showed a more accurate color distribution in Test-S300, and it yielded results similar to DM-water in Test-U90, while being closer to the reference image in terms of color histogram and saliency feature detection results.
 In summary, WaterFormer exhibited outstanding performance in the UIE task and proves to be the most beneficial in related applications.
 4.5 Ablation Study
-The proposed WaterFormer incorporates three major components: the CRB, CFB, and loss of combining ${{L}_{chroma}}$ and ${{L}_{Sobel}}$.
-To demonstrate the effectiveness of each component, an ablation study was performed on the baseline framework with various combinations of these components.
-Tab.
-3 shows the statistical results on the Test-S300 dataset, where the full model ($V_{5}$) outperformed other variants in terms of SSIM and PSNR scores.
-Variants CRB CFB ${{L}_{chroma}}$ ${{L}_{Sobel}}$ SSIM↑ PSNR↑
-Base w/o w/o w/o w/o 0.872 26.64
-${{V}_{1}}$ w/o w/o w/o 0.876 27.09
-${{V}_{2}}$ w/o w/o 0.877 27.26
-${{V}_{3}}$ w/o 0.891 28.47
-${{V}_{4}}$ w/o 0.901 28.56
-${{V}_{5}}$ 0.907 29.40
-Table 3: Quantitative analysis of the WaterFormer ablation experiments. The red font indicates the best performance.
-Fig.
-10 displays the visual results of various variants of WaterFormer.
-The Base model has basic UIE capabilities but retained significant shortcomings in addressing the color degradation and edge artifacts.
-The $V_{1}$ model, which incorporates the CRB module, could more delicately restore colors.
-However, due to insufficient feature fusion, the results of the $V_{1}$ model remained lacking.
-Therefore, by introducing the CFB module, the predictions of the $V_{2}$ model more closely aligned with the true color distribution.
-The addition of ${{L}_{chroma}}$ in the $V_{3}$ model further enhanced the consistency between inference results and chromaticity of the real images.
-However, $V_{1}$, $V_{2}$, and $V_{3}$ continued to exhibit noticeable edge artifacts, and this problem was addressed by the inclusion of ${{L}_{Sobel}}$, as demonstrated in the results of the $V_{4}$ inference.
-Finally, by combining all components, our WaterFormer ($V_{5}$) yielded the closest results to the real images.
+The proposed WaterFormer incorporates four main components: the network backbone, CRB, CFB, and the loss function formed by ${{L}_{\ell_{1}}}$, ${{L}_{chroma}}$ and ${{L}_{Sobel}}$.
+To validate the effectiveness of each component, a series of ablation studies were conducted on the baseline framework with different combinations of these components.
+The Test-S300 was used for the ablation experiments, with each image resized to $256\times 256$.
+The experiments were divided into four parts: first, WaterFormer was broken down into various variants, each composed of different components, and tested on Test-S300; second, the superiority of the CRB design was further explored, and different attention modules were tested as replacements for CRB; third, the CFB fusion mechanism was visualized, with pixel-wise addition used as a comparison; fourth, other optimizations and experimental parameters of WaterFormer were tested by ablation.
 […]
 Fig. 10: Qualitative analysis of the WaterFormer ablation experiments. (a) Input, (b) Base, (c) $V_{1}$, (d) $V_{2}$, (e) $V_{3}$, (f) $V_{4}$, (g) $V_{5}$, (h) GT.
-In addition, we will demonstrate the effectiveness of other improvements on WaterFormer in Tab.
-4, including replacing the ReLU with the FReLU [39] in the MLP structure and replacing the global residual reconstruction layer with the soft reconstruction layer that was proposed in [13] and optimized according to Eq.
+Variants CRB CFB ${{L}_{chroma}}$ ${{L}_{Sobel}}$ SSIM↑ PSNR↑
+Base w/o w/o w/o w/o 0.876 27.32
+${{V}_{1}}$ w/o w/o w/o 0.890 28.50
+${{V}_{2}}$ w/o w/o 0.892 28.87
+${{V}_{3}}$ w/o 0.907 30.35
+${{V}_{4}}$ w/o 0.913 30.46
+${{V}_{5}}$ 0.917 31.18
+Table 4: Quantitative analysis of the WaterFormer ablation experiments. The bold font indicates the best performance.
+Variants Evaluation.
+Fig.
+10 and Fig.
+4 display the statistical results and visualization results on the Test-S300.
+The base model was composed only of the backbone and was trained using the ${{L}_{\ell_{1}}}$.
+It had basic UIE capabilities but still retained significant shortcomings in addressing color degradation and edge artifacts.
+The $V_{1}$ model introduced the CRB, which enabled it to recover color more accurately.
+However, due to insufficient integration of different features, the results of $V_{1}$ still exhibited a large amount of color artifacts.
+Therefore, by introducing the CFB, the predictions of the $V_{2}$ model exhibited fewer color artifacts and a more realistic color distribution.
+The addition of ${{L}_{chroma}}$ to the $V_{3}$ model further improved the consistency between inference results and the chromaticity of real images.
+However, $V_{1}$, $V_{2}$, and $V_{3}$ continued to exhibit noticeable edge artifacts, such as the white rim of the bucket, and this problem was addressed by including ${{L}_{Sobel}}$, as demonstrated in the results of the inference $V_{4}$.
+Finally, by combining all components, our WaterFormer ($V_{5}$) yielded the closest results to the real images with the highest SSIM and PSNR scores.
+CRB Evaluation.
+The design of CRB is based on the mechanism of self-attention between channels, as proposed by  [34] and  [35].
+CRB extends this mechanism and applies it effectively to the UIE task.
+To visualize the effect of CRB,  11 shows the output of the feature map of the first 2-block CRB in the $V_{1}$ variant.
+For comparison, the feature map output of the first 2-block DehazeFormer Block in the Base model is included.
+We used maximum projection and the jet color map to plot the heatmap, where blue represents lower attention from the network, and red represents higher attention.
+In Test-S300, we selected two representative types of water for validation.
+The first row represents clear water, with only color degradation, but clear foreground and background.
+The second row represents turbid water, with both color degradation and severe occlusion.
+As seen in Fig.
+11, the DehazeFormer Block focuses on the foreground, and its performance decreases significantly when most of the foreground is occluded.
+In contrast, CRB can focus on global color correction, and its heatmap more accurately reflects the distribution of color bias.
+This shows that the DehazeFormer Block and CRB handle different subtasks in the UIE task, with CRB assisting the DehazeFormer Block in feature extraction while performing color restoration.
+Furthermore, to demonstrate the higher performance brought about by CRB, we replaced MTDA in  [34] and MCC in [35] with CRB.
+The results are shown in Tab.
+5.
+Additionally, RLN in the CRB significantly improved color restoration, and for comparison, BatchNorm and LayerNorm were included in the ablation study.
+To further highlight the benefits of CRB and channel self-attention, we also included popular attention mechanisms, SE Block [44] and CBAM [45], in the ablation experiment.
+[…]
+Fig. 11: Visualization of the CRB working mechanism. (a) Input, (b) result of DehazeFormer Block, (c) result of CRB, (d) GT.
+Variants SSIM↑ PSNR↑
+WaterFormer 0.917 31.18
+CRB→MTDA [34] 0.910 30.01
+CRB→MCC [35] 0.912 30.43
+CRB→SE Block [44] 0.909 30.00
+CRB→CBAM [45] 0.907 29.89
+RLN→BatchNorm 0.916 31.05
+RLN→LayerNorm 0.915 30.85
+Table 5: Quantitative analysis of CRB ablation experiments. The bold font indicates the best performance.
+CFB Evaluation.
+To better validate the effectiveness of the CFB, its fusion mechanism was visualized and compared with pixel-wise addition, as shown in Fig.
+12.
+The first CFB example is displayed, where its two input feature maps come from the image embedding layer (composed of a convolutional layer with a 3x3 kernel size) and the 2-block CRB.
+We also selected clear and turbid water types for the experiment in Test-S300, which are presented in the first and second rows, respectively.
+In Fig.
+12, the image embedding layer can only extract shallow features, and its heatmap reflects mainly the pixel distribution of the input.
+On the other hand, as described in the previous subsection, the heatmap generated by CRB displays the distribution of color bias information.
+From the heatmap, it can be seen that if the two feature maps are added pixel-wise, many details are lost, and the WaterFormer inference results are therefore affected.
+However, after fusion using the CFB, the results for both clear and turbid water are presented with more distinct layers and better detail preservation (as indicated by the purple arrow).
+[…]
+Fig. 12: Visualization of the CFB fusion mechanism. (a) Input, (b) result of image embedding layer, (c) result of CRB, (d) result of CFB, (e) result of addition, (f) GT.
+Other Evaluation.
+The effectiveness of other improvements in WaterFormer was demonstrated in Tab.
+6, including the replacement of the ReLU with the FReLU [33] in the MLP structure and the replacement of the global residual reconstruction layer with the soft reconstruction layer proposed in [14] and optimized according to Eq.
 3.
-References [13, 52] also proposed the SK Fusion structure to fuse features from different stages.
-Unfortunately, we found that the SK Fusion structure was not suitable for feature fusion in underwater images, possibly because SK Fusion does not effectively utilize the global information among different feature channels in underwater images.
-We present the ablation experiments on SK Fusion in Tab.
-4.
+The study on the size of the sliding window in ${{L}_{chroma}}$ can be found in Tab.
+7.
+The study on the scaling coefficients of each loss term in the loss function can be found in Tab.
+8.
 Variants SSIM↑ PSNR↑
-WaterFormer 0.907 29.40
-FReLU [39]→ReLU 0.905 29.40
-UW Soft Recon→Recon 0.904 29.00
-UW Soft Recon→Soft Recon [13] 0.904 28.93
-CFB→SK fuison [13, 52] 0.903 28.88
-Table 4: Quantitative analysis of the ablation experiments on other improvements in WaterFormer. The red font indicates the best performance.
+WaterFormer 0.917 31.18
+FReLU [33]→ReLU 0.915 30.82
+UW Soft Recon→Recon 0.911 30.26
+UW Soft Recon→Soft Recon [14] 0.916 30.79
+Table 6: Quantitative analysis of the ablation experiments on other improvements in WaterFormer. The bold font indicates the best performance.
+Sliding window size SSIM↑ PSNR↑
+$11\times 1$ 11 0.916 31.12
+$13\times 13$ 0.916 31.04
+$15\times 15$ 0.917 31.18
+Table 7: Quantitative analysis of the ablation experiments on size of the sliding window in ${{L}_{chroma}}$. The bold font indicates the best performance.
+${\lambda}_{1}$ ${\lambda}_{2}$ ${\lambda}_{3}$ SSIM↑ PSNR↑
+1 1 1 0.912 30.53
+2 1 1 0.909 30.03
+1 2 1 0.915 30.88
+1 1 2 0.917 31.18
+2 2 1 0.912 30.48
+2 1 2 0.912 30.84
+1 2 2 0.916 31.29
+2 2 2 0.911 30.20
+Table 8: Quantitative analysis of the ablation experiments on scaling coefficients of each loss term in the loss function. The bold font indicates the best performance.
 5 Conclusion
-In this work, we introduced WaterFormer, a versatile method to enhance the quality of underwater images.
-To tackle challenges such as haziness and color shifts, WaterFormer leverages the Vision Transformer architecture and comprises three key components: the DehazeFormer Block, Color Restoration Block (CRB), and Channel Fusion Block (CFB).
-The DehazeFormer Block transforms hazy features and extract deep-level features.
+In this work, we introduce WaterFormer, a versatile method to enhance the quality of underwater images.
+To address challenges such as haziness and color shifts, WaterFormer leverages the ViT architecture and comprises three key components: the DehazeFormer Block, the Color Restoration Block (CRB) and the Channel Fusion Block (CFB).
+The DehazeFormer Block transforms hazy features and extracts deep-level features.
 The CRB enhances colors through channel-level transpose matrices.
@@
 We also incorporated the Chromatic Consistency Loss and Sobel Color Loss to ensure color fidelity and preserve fine details.
-The quantitative evaluations, qualitative evaluations, and ablation experiments show that our proposed network performs significantly better than various methods.
-Nevertheless, this study has several limitations.
+Quantitative evaluations, qualitative evaluations, and ablation experiments show that our proposed network performs significantly better than various methods.
+However, this study has several limitations.
 First, our work requires a substantial amount of paired data for training, which poses a significant challenge for practical applications.
-Second, we did not consider the effects of noise such as environmental noise and camera noise during the image processing.
+Second, we did not consider the effects of noise, such as environmental noise and camera noise, during image processing.
 Large-scale datasets that cover and fully consider complex types of image noise will be essential to train a more robust and generalizable model.
 In our subsequent work, we will consider incorporating a denoising module into the network to reduce the impact of noise and enhance the robustness of the image correction results.
-In addition, we will use a contrastive learning strategy to reduce the network dependency on data and improve the generalization performance.
+In addition, we will use a contrastive learning strategy to reduce the dependency of the network on data and improve generalization performance.
 The denoising process will be integrated into the end-to-end network as a module instead of a separate preprocessing step.